OpenStudy (anonymous):

need help! The Equation of the sides of a quadrilateral ABCD are as follows AB: 4y+x=20 BC: x=0 CD:y+2x=2 AD: 3y=4x-4 Find the coordinates of A, B, C and D a=by a graphical method

OpenStudy (binary3i):

solve all of them, then see such points which are only on two lines not more than that.

OpenStudy (callisto):

Solve AB: 4y+x=20 BC: x=0 To get B Solve BC: x=0 CD:y+2x=2 To get C Solve CD:y+2x=2 AD: 3y=4x-4 To get D Solve AB: 4y+x=20 AD: 3y=4x-4 To get A

OpenStudy (anonymous):

thank you!

OpenStudy (callisto):

Welcome~